# Trial and Trust: Addressing Byzantine Attacks with Comprehensive Defense Strategy

### v1

@@ Appendix C General Inequalities and Lemmas
$\begin{split}\quad\varphi(y)\quad&\geqslant\quad\varphi(x)+\left\langle\nabla\varphi(x),y-x\right\rangle\\ 0\quad&\geqslant\quad\left\langle\varphi(x)-\varphi(y),y-x\right\rangle\end{split}$ (Conv)
@@
To establish the uniform convergence of $\left\|\nabla f_{1}(x)-\nabla\hat{f}(x)\right\|_{\infty}^{2}$, we employ Theorem 5 from (Shalev-Shwartz et al., 2009). This theorem provides a bound on the $\ell_{\infty}$-covering number of the function class =$\mathcal{F}=\{\xi\mapsto\nabla f_{1}(x;\xi)\mid x\in\mathcal{X}\}$. Given that $\mathcal{X}$ resides within an $\ell_{2}$-sphere, let us define it bound by $S$, the covering number for $\mathcal{X}$ using the Euclidean metric $d_{2}(x_{i},x_{j})=\|x_{i}-x_{j}\|_{2}$ is constrained as follows for $d>3$:
@@
Following this analysis, we derive an estimation consistent with the findings in (Shalev-Shwartz et al., 2009):
@@
[…]
@@
$-\gamma\left\langle\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$ $=$ $\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle.$
@@
$\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$ (${\leqslant}$ $\frac{\gamma}{2}\left\|\nabla f(x)-\nabla\hat{f}(x)\right\|^{2}+\frac{\gamma}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{1}(x)\right\|^{2}$
(${\leqslant}$ $\frac{\gamma}{2}\left(\left\|\nabla f(x)-\nabla f_{1}(x)\right\|^{2}+\left\|\nabla f_{1}(x)-\nabla\hat{f}(x)\right\|^{2}\right)$
$+\frac{\gamma}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|^{2}$
@@
$-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$ ${=}$ $-\frac{\gamma}{2}\|\nabla f(x)\|^{2}-\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|^{2}$
@@
(${\leqslant}$ $-\frac{\gamma}{2}\|\nabla f(x)\|^{2}-\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|^{2}$
@@ Appendix D Proofs of BANT
Under Assumptions 2.1, 2.2(b), 2.5, the following holds for the iteration of Algorithm 1:
@@
(${\leqslant}$ $\hat{f}(x^{t})+\frac{\beta}{n}\sum\limits_{i\in\mathcal{G}}\Biggl{[}\hat{f}(x^{t})-\gamma\left\langle\nabla\hat{f}(x^{t}),g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\|g_{i}^{t}\|^{2}-\hat{f}(x^{t})\Biggr{]}$
@@
Theorem 3.1. Under Assumptions 2.1, 2.2(b), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{12}$, 2.5, for solving the problem described in the equation (1) after $T$ iterations of Algorithm 1 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
@@
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\frac{\gamma\beta G}{n}\zeta(N)-\frac{\gamma\beta G}{2n}\left\|\nabla f(x^{t})\right\|^{2}+\frac{3\gamma\beta G}{2n}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}\beta}{2n}\left(\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}+\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)+\frac{3L\gamma^{2}\beta G}{2n}\left\|\nabla f(x^{t})\right\|^{2}$
@@
$\frac{1}{T}\sum\limits_{i=1}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$ $=$ $\mathcal{O}\Biggl{(}\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}{\beta GT}\quad+\frac{\sigma\cdot\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}}{\sqrt{\beta GT}}+\delta_{1}+\zeta(N)\Biggr{)};$
@@
We proceed estimation, analogical to Lemma 4 from (Stich, 2019). Using the result of Theorem 3.1, we choose the appropriate $\gamma\leqslant\min\left\{\frac{1}{13L},\frac{\sqrt{2\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]n}}{\sigma\sqrt{3LG\beta T}}\right\}$. In that way, we obtain
@@
Under Assumptions 2.1, 2.2(a), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{12}$, 2.5, for solving the problem described in the equation (1) after $T$ iterations of Algorithm 1 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
@@
(${\geqslant}$ $\frac{1}{2}\mathbb{E}\left\|\nabla f(x^{t})+\nabla\hat{f}(x^{t})-\nabla f(x^{t})\right\|^{2}-\mathbb{E}\left\|\nabla\hat{f}(x^{t})-\nabla f(x^{t})\right\|^{2}$
@@
$\mathbb{E}\left[\hat{f}(x^{t})-\hat{f}(\hat{x}^{*})\right]$$\leqslant$ $\left(1-\gamma^{\prime}\mu\right)^{t}\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]+\underbrace{\sum\limits_{i=0}^{t-1}\left(1-\gamma^{\prime}\mu\right)^{i}}_{\leqslant\frac{1}{\gamma^{\prime}\mu}}\left[8\gamma^{\prime}\delta_{1}+9L\gamma\gamma^{\prime}\sigma^{2}+9\gamma^{\prime}\zeta(N)\right].$
@@
In Theorem D.2, we obtain classic result for SGD. We use Lemma 2 from (Stich, 2019) and appropriate special tunings of $\gamma$:
@@ Appendix E Proofs of AutoBANT
Our goal is to learn how to circumvent this limitation. To achieve this, we tackle an additional subproblem related to weight assignment. This step represents a key distinction in our theoretical analysis. We can assert that we are solving this weight distribution minimization subproblem with a certain error margin $\delta$, which will be reflected in the final convergence results. Next, we present the theorem discussed in the main part of the paper, along with its complete proof.
Theorem 3.3. Under Assumptions 2.1, 2.2(c), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{12}$, 2.5, for solving the problem described in the equation (1) after $T$ iterations of Algorithm 2 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{G}\sigma^{2}+4\zeta(N)+\frac{4\delta}{\gamma}.$
@@
$\hat{f}(x^{t+1})$$\leqslant$ $\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[x^{t}-\gamma\sum_{i=1}^{n}\omega_{i}g_{i}^{t}\right]+\delta$
$\leqslant$ $\hat{f}\left[x^{t}-\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right]+\delta$
(${\leqslant}$ $\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
@@
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
@@
$+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\gamma\zeta(N)-\frac{\gamma}{2}\left\|\nabla f(x^{t})\right\|^{2}+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}(\nabla f(x^{t})-\nabla f_{i}(x^{t}))\right\|^{2}+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}(\nabla f_{i}(x^{t})-g_{i}^{t})\right\|^{2}$
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f(x^{t})\right\|^{2}+\delta.$
@@
$\mathbb{E}\hat{f}(x^{t+1})$ (${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\gamma\zeta(N)-\frac{\gamma}{2}\left\|\nabla f(x^{t})\right\|^{2}+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
@@
$+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
@@
$+\frac{3L\gamma^{2}}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\frac{\sigma^{2}}{G}\right)+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
@@
$+\frac{2\gamma}{2}(1+3L\gamma)\delta_{1}+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma\zeta(N)+\delta.$
@@
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{G}\sigma^{2}+4\zeta(N)+\frac{4\delta}{\gamma}.$
@@ Appendix F Scaled methods
This update scheme is satisfied by Adam-based methods with $(H^{t})^{2}=\text{diag}(g^{t}\odot g^{t})$ and by AdaHessian (Yao et al., 2021) with $(H^{t})^{2}=\text{diag}(z^{t}\odot\nabla^{2}f(x^{t}))^{2}$, where $\odot$ denotes the component-wise product between two vectors, and $z^{t}$ are from Rademacher distribution, i.e. all components from vector are independent and equal to ±$\pm 1$ with probability $\nicefrac{{1}}{{2}}$.
@@
For example, such a rule is extended in OASIS (Jahani et al., 2022) with $\beta_{t}\equiv\beta$ and $H^{t}=\text{diag}(z^{t}\odot\nabla^{2}f(x^{t}))$. We also note additional details in the construction of a positively defined preconditioner. We can also alternatively define $\hat{P}^{t}$ as $(\hat{P})_{ii}=|P|_{ii}+e$. Most importantly, both of these approaches construct diagonal matrices with positive elements. We introduce the crucial for our analysis assumption.
@@
The correct proof of this statement, as well as a more detailed description of the diagonal preconditioner, is provided in (Sadiev et al., 2024). We mention that for AdaHessian and OASIS preconditioners, $\Gamma=\sqrt{d}L$, and for Adam and RMSProp, under the condition $\|\nabla f(x)\|\leqslant M,\Gamma=M$. Thus, having constructed a diagonal preconditioner, we can proceed to the analysis of scaled methods.

### v5

@@ Appendix C General Inequalities and Lemmas
$\begin{split}\quad\varphi(y)\quad&\geqslant\quad\varphi(x)+\left\langle\nabla\varphi(x),y-x\right\rangle\\ 0\quad&\geqslant\quad\left\langle\nabla\varphi(x)-\nabla\varphi(y),y-x\right\rangle\end{split}$ (Conv)
@@
To establish the uniform convergence of $\left\|\nabla f_{1}(x)-\nabla\hat{f}(x)\right\|_{\infty}^{2}$, we employ Theorem 5 from [shalev2009stochastic]. This theorem provides a bound on the $\ell_{\infty}$-covering number of the function class $\mathcal{F}=\{\xi\mapsto\nabla f_{1}(x;\xi)\mid x\in\mathcal{X}\}$. Given that $\mathcal{X}$ resides within an $\ell_{2}$-sphere, let us define it bound by $S$, the covering number for $\mathcal{X}$ using the Euclidean metric $d_{2}(x_{i},x_{j})=\|x_{i}-x_{j}\|_{2}$ is constrained as follows for $d>3$:
@@
Following this analysis, we derive an estimation consistent with the findings in [shalev2009stochastic]:
@@
$\zeta(N)\overset{\text{def}}{=}\widetilde{\mathcal{O}}\left(\frac{1}{N}\right),$
@@
$-\gamma\left\langle\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$ $=$ $\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$
$-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle.$
@@
$\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle$
[…]
@@
[…]
@@
[…]
@@ Appendix D Proofs of Bant
Under Assumptions 3.1, 3.2(b), 3.5, the following holds for the iteration of Algorithm 1:
@@
[…]
@@
Theorem 4.1. Under Assumptions 3.1, 3.2(b), 3.3, 3.4 with $\delta_{2}\leqslant\frac{1}{12}$, 3.5, for solving the problem described in the equation (1) after $T$ iterations of Algorithm 1 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
@@
[…]
$+\frac{3\gamma\beta G}{2n}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}\beta}{2n}\left(\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}+\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)$
$+\frac{3L\gamma^{2}\beta G}{2n}\left\|\nabla f(x^{t})\right\|^{2}$
@@
$\frac{1}{T}\sum\limits_{i=1}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}=\mathcal{O}\Biggl(\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}{\beta GT}+\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}}{\sqrt{\beta GT}}+\delta_{1}+\zeta(N)\Biggr).$
@@
We proceed estimation, analogical to Lemma 4 from [stich2019unified]. Using the result of Theorem 4.1, we choose the appropriate $\gamma\leqslant\min\left\{\frac{1}{13L},\frac{\sqrt{2\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]n}}{\sigma\sqrt{3LG\beta T}}\right\}$. In that way, we obtain
@@
Under Assumptions 3.1, 3.2(a), 3.3, 3.4 with $\delta_{2}\leqslant\frac{1}{12}$, 3.5, for solving the problem described in the equation (1) after $T$ iterations of Algorithm 1 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
@@
[…]
@@
$\mathbb{E}\left[\hat{f}(x^{t})-\hat{f}(\hat{x}^{*})\right]$ $\leqslant$ $\left(1-\gamma^{\prime}\mu\right)^{t}\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]$
$+\underbrace{\sum\limits_{i=0}^{t-1}\left(1-\gamma^{\prime}\mu\right)^{i}}_{\leqslant\frac{1}{\gamma^{\prime}\mu}}\left[8\gamma^{\prime}\delta_{1}+9L\gamma\gamma^{\prime}\sigma^{2}+9\gamma^{\prime}\zeta(N)\right],$
@@
In Theorem D.2, we obtain classical result for SGD. We use Lemma 2 from [stich2019unified] and appropriate special tunings of $\gamma$:
@@ Appendix E Proofs of AutoBant
Our goal is to learn how to circumvent this limitation. To achieve this, we tackle an additional subproblem related to weight assignment. This step represents a key distinction in our theoretical analysis. We assert that we can solve this weight distribution minimization subproblem to any predefined accuracy. Next, we present the theorem discussed in the main part of the paper, along with its complete proof.
Theorem 4.3. Under Assumptions 3.1, 3.2(c), 3.3, 3.4 with $\delta_{2}\leqslant\frac{1}{12}$, 3.5, for solving the problem described in the equation (1) after $T$ iterations of Algorithm 2 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$ $\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{G}\sigma^{2}+4\zeta(N).$
@@
$\hat{f}(x^{t+1})$ $=$ $\min_{\omega\in\Delta_{1}^{n}}\hat{f}\left[x^{t}-\gamma\sum_{i=1}^{n}\omega_{i}g_{i}^{t}\right]$
$\leqslant$ $\hat{f}\left[x^{t}-\frac{\gamma}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right]$
[…]
@@
$\mathbb{E}\hat{f}(x^{t+1})$ $\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\left\langle\nabla\hat{f}(x^{t}),\frac{\gamma}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$+\frac{3\gamma}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
[…]
$+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}(\nabla f(x^{t})-\nabla f_{i}(x^{t}))\right\|^{2}$
$+\frac{3L\gamma^{2}}{2}\mathbb{E}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}(\nabla f_{i}(x^{t})-g_{i}^{t})\right\|^{2}+\frac{3L\gamma^{2}}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f(x^{t})\right\|^{2}.$
@@
[…]
@@
$+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}$
@@
$+\frac{3L\gamma^{2}}{2}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\frac{\sigma^{2}}{G}\right)+\frac{3L\gamma^{2}}{2}\left\|\nabla f(x^{t})\right\|^{2}$
@@
$+\frac{2\gamma}{2}(1+3L\gamma)\delta_{1}+\frac{3L\gamma^{2}}{2G}\sigma^{2}+\gamma\zeta(N).$
@@
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$ $\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{G}\sigma^{2}+4\zeta(N).$
@@ Appendix F Scaled methods
This update scheme is satisfied by Adam-based methods with $(H^{t})^{2}=\text{diag}(g^{t}\odot g^{t})$ and by AdaHessian [yao2021adahessian] with $(H^{t})^{2}=\text{diag}(z^{t}\odot\nabla^{2}f(x^{t}))^{2}$, where $\odot$ denotes the component-wise product between two vectors, and $z^{t}$ are from Rademacher distribution, i.e. all components from vector are independent and equal to $\pm 1$ with probability $\nicefrac{{1}}{{2}}$.
@@
For example, such a rule is extended in OASIS [jahani2021doubly] with $\beta_{t}\equiv\beta$ and $H^{t}=\text{diag}(z^{t}\odot\nabla^{2}f(x^{t}))$. We also note additional details in the construction of a positively defined preconditioner. We can also alternatively define $\hat{P}^{t}$ as $(\hat{P})_{ii}=|P|_{ii}+e$. Most importantly, both of these approaches construct diagonal matrices with positive elements. We introduce the crucial for our analysis assumption.
@@
The correct proof of this statement, as well as a more detailed description of the diagonal preconditioner, is provided in [sadiev2024stochastic]. We mention that for AdaHessian and OASIS preconditioners, $\Gamma=\sqrt{d}L$, and for Adam and RMSProp, under the condition $\|\nabla f(x)\|\leqslant M,\Gamma=M$. Thus, having constructed a diagonal preconditioner, we can proceed to the analysis of scaled methods.